OpenStudy (anonymous):

Estimate the sum or difference. 24 1/10 + 5 9/11 A 19 B 18 C 17

OpenStudy (studygurl14):

Estimate: \(\large 24\frac{1}{10}\rightarrow 24\) \(\large 5\frac{9}{11}\rightarrow 6\) 24 + 6 = ?